A potential difference Δ? exists between the inner and outer surfaces of the membrane of a cell. The inner surface is negative relative to the outer surface. If 1.10×10−20 J of work is required to eject a positive sodium ion (Na+) from the interior of the cell, what is the magnitude of the potential difference (in millivolts) between the inner and outer surfaces of the cell?
Answer #1

answer) here we have the formula

w=q\DeltaV

\Deltav=w/q=1.10*10-20J/1.60*10-19C=0.6875V=687.5 mV

so the answer is 688mV or 687.5mV.